Required information

Problem 11-4A Analysis of changes in stockholders' equity accounts LO C3, P2, P3

[The following information applies to the questions displayed below.]

The equity sections from Atticus Group’s 2016 and 2017 year-end balance sheets follow.

Stockholders’ Equity (December 31, 2016)
Common stock—$5 par value, 100,000 shares
authorized, 40,000 shares issued and outstanding
$ 200,000
Paid-in capital in excess of par value, common stock 160,000
Retained earnings 320,000
Total stockholders’ equity $ 680,000
Stockholders’ Equity (December 31, 2017)
Common stock—$5 par value, 100,000 shares
authorized, 47,000 shares issued, 5,000 shares in treasury
$ 235,000
Paid-in capital in excess of par value, common stock 195,000
Retained earnings ($30,000 restricted by treasury stock) 400,000
830,000
Less cost of treasury stock (30,000 )
Total stockholders’ equity $ 800,000


The following transactions and events affected its equity during year 2017.

Jan. 5 Declared a $0.60 per share cash dividend, date of record January 10.
Mar. 20 Purchased treasury stock for cash.
Apr. 5 Declared a $0.60 per share cash dividend, date of record April 10.
July 5 Declared a $0.60 per share cash dividend, date of record July 10.
July 31 Declared a 20% stock dividend when the stock’s market value was $10 per share.
Aug. 14 Issued the stock dividend that was declared on July 31.
Oct. 5 Declared a $0.60 per share cash dividend, date of record October 10.

Problem 11-4A Part 1

Required:

1. How many common shares are outstanding on each cash dividend date?
Jan. 5 Apr. 5 July 5 Oct. 5 Outstanding common shares

2. What is the total dollar amount for each of the four cash dividends?

Jan. 5 Apr. 5 July 5 Oct. 5 Cash dividend amounts

3. What is the amount of the capitalization of retained earnings for the stock dividend?

Capitalization amount Capitalization amount

4. What is the per share cost of the treasury stock purchased?

Capitalization amount Capitalization amount

5. How much net income did the company earn during year 2017?

Net income Net income L
Answer #1

1.

Jan. 5 Apr. 5 July 5 Oct. 5
Outstanding common shares 40000 35000 35000 42000
(40000 - 5000) (1.2 x 35000)

2.

Jan. 5 Apr. 5 July 5 Oct. 5
Cash dividend amounts 24000 21000 21000 25200
(40000 x $0.60) (35000 x $0.60) (35000 x $0.60) (42000 x $0.60)

3. Capitalization amount = 20% x 35000 x $10 = $70000

4. Cost per share = $30000/5000 = $6

5. Net income: $241200

RE(beg) + NI - Cash dividends - Stock divided = RE(end)

$320000 + NI - $91200 - $70000 = $400000

$158800 + NI = $400000

NI = $400000 - $158800 = $241200